OpenStudy (anonymous):

How do i take the derivative of 2x^2 - 200x + 10000

OpenStudy (anonymous):

just bring down the exponent which is 4x - 200

OpenStudy (anonymous):

thats all you did

OpenStudy (unklerhaukus):

the derivative with respect to x?

OpenStudy (anonymous):

so the x on 200x disappears?

OpenStudy (anonymous):

Apply Power rule!

OpenStudy (anonymous):

what about the 10000?

OpenStudy (anonymous):

(c)' = 0

OpenStudy (anonymous):

the 10000 doesn't do anything

OpenStudy (anonymous):

Yeh, power rule mang. The 200x is implied to be 200x^1. 1 * 200 = 200. 1 - 1 = 0 and x^0 = 1. So 200 * 1.

OpenStudy (anonymous):

it equals 0

OpenStudy (unklerhaukus):

\[y(x)=2x^2 - 200x + 10000\] \[\frac{\text dy(x)}{\text dx}=\frac{\text d}{\text dx}\left(2x^2 - 200x + 10000\right)\] \[=\frac{\text d}{\text dx}(2x^2) - \frac{\text d}{\text dx}(200x) + \frac{\text d}{\text dx}(10000)\]\[=(2\times2x^{2-1}) - (1\times200x^{1-1}) +(0\times10000x^{0-1})\]